# SynFER: Towards Boosting Facial Expression Recognition with Synthetic Data

### v2

@@ Abstract
Facial expression datasets remain limited in scale due to privacy concerns, the subjectivity of annotations, and the labor-intensive nature of data collection. This limitation poses a significant challenge for developing modern deep learning-based facial expression analysis models, particularly foundation models, that rely on large-scale data for optimal performance. To tackle the overarching and complex challenge, we introduce SynFER (Synthesis of Facial Expressions with Refined Control), a novel framework for synthesizing facial expression image data based on high-level textual descriptions as well as more fine-grained and precise control through facial action units. To ensure the quality and reliability of the synthetic data, we propose a semantic guidance technique to steer the generation process and a pseudo-label generator to help rectify the facial expression labels for the synthetic images. To demonstrate the generation fidelity and the effectiveness of the synthetic data from SynFER, we conduct extensive experiments on representation learning using both synthetic data and real-world data. Experiment results validate the efficacy of the proposed approach and the synthetic data. Notably, our approach achieves a 67.23% classification accuracy on AffectNet when training solely with synthetic data equivalent to the AffectNet training set size, which increases to 69.84% when scaling up to five times the original size. Our code will be made publicly available.
@@ 1 Introduction
Facial Expression Recognition (FER) is at the forefront of advancing AI’s ability to interpret human emotions, opening new frontiers for various human-centered applications. From automatic emotion detection to early interventions in mental health Ringeval et al. (2019), accurate pain assessment Huang et al. (2024), and enhancing human-computer interaction Abdat et al. (2011), the potential impact of FER systems is profound Moin et al. (2023); Sajjad et al. (2023); Zhu & Luo (2023). In recent years, learning-based FER models have gained significant traction due to their promising performances Li & Deng (2020); Zhang et al. (2021); Farzaneh & Qi (2021). However, despite recent advancements in network architectures and learning methodologies, the progress of existing FER models has been hindered by the inadequate scale and quality of available training data, underscoring the need to expand datasets with high-quality data to push the boundaries of FER capabilities.
@@
Existing FER datasets, such as CK+ (953 sequences) Lucey et al. (2010), FER-2013 (30,000 48$\times$48 images) Barsoum et al. (2016), RAF-DB (29,672 images) Li et al. (2017), AFEW (113,355 images) Dhall et al. (2017), and SFEW (1,766 images) Dhall et al. (2011), are small compared to popular image datasets for general image processing (e.g., ImageNet Deng et al. (2009) with 1.4 million images and Laion Schuhmann et al. (2022) with billion-level data). While AffectNet Mollahosseini et al. (2017) compiles a large number of facial images from the web, it still suffers from vital drawbacks. A considerable portion of AffectNet’s images are low-quality, and its annotations often contain incorrect labels, which impairs the training process of FER models Le et al. (2023); Yan et al. (2022). Consequently, the absence of high-quality and large-scale FER datasets has delayed the development of FER foundation models. However, collecting a large-scale FER dataset with high-quality facial images and meticulous annotations is almost an unrealistic endeavor due to substantial financial and time costs, ethical concerns around facial data collection, and limited resources for large-scale acquisition. Additionally, the subjective interpretation of facial expressions results in inconsistent labeling by annotators, which exacerbates variability and hinders the creation of reliable datasets.
To address the challenges in developing FER models, we turn to synthesizing high-quality facial expression images paired with reliable labels. This approach draws inspiration from successful strategies employed to expand annotated datasets for other computer vision tasks, such as semantic segmentation Baranchuk et al. (2021); Chen et al. (2019); Li et al. (2022a) and depth estimation Atapour-Abarghouei & Breckon (2018); Cheng et al. (2020); Guizilini et al. (2022). These advances leverage powerful generative models such as Stable Diffusion Rombach et al. (2022a) and DALL-E Betker et al. (2023), which capture intricate natural image patterns. By tapping into these models, researchers have generated realistic images with their corresponding annotations, thereby boosting model performance. However, applying diffusion models to synthesize facial expression images with reliable FER labels presents two major challenges. (1) the training sets used by these generative models often lack diverse facial expression data, limiting their ability to produce images that capture subtle and nuanced emotional semantics; and (2) prior approaches to generate annotations for synthetic images focused on tangible attributes such as pixel-wise layouts, or depth maps. In contrast, facial expressions convey abstract and subjective emotions, making the generation of precise and reliable expression labels much more complex. To the best of our knowledge, none of the existing methods can simultaneously conduct fine-grained control for facial expression generation and generate robust categorical facial expression labels for face images.
@@
In this paper, we present SynFER, the first framework capable of synthesizing unlimited, diverse and realistic facial expression images paired with reliable expression labels, to drive advancements in FER models. To address the shortcomings of existing FER datasets, which often lack expression-related text paired with facial images, we introduce FEText, a unique hybrid dataset created by curating and filtering data from existing FER and high-quality face datasets. This vision-language dataset serves as the foundation for training our generative model to synthesize facial expression data. To ensure fine-grained control and faithful generation of facial expression images, we inject facial action unit (FAU) information and semantic guidance from external pre-trained FER models. Building upon this, we propose FERAnno, the first diffusion-based label calibrator for FER, which automatically generates reliable annotations for the synthesized images. Together, these innovations position SynFER as a powerful tool for producing large-scale, high-quality facial expression data, offering a significant resource for the development of FER models.
@@
We investigate the effectiveness of the synthetic data across various learning paradigms, demonstrating consistent and modest improvement in model performance. As shown in Fig. 1(c), training with the synthetic data yields significant performance boosts across various learning paradigms. Notably, pre-training on the synthetic data (Fig. 1(b)) with MoCo v3 Chen et al. (2021) yields a significant performance boost of +2.90% on AffectNet, surpassing real-world data pre-training. In supervised learning, SynFER improves accuracy by +1.55% for the state-of-the-art FER model, POSTER++ Mao et al. (2024), on AffectNet. We further explore performance scaling of the synthetic data, revealing further gains as dataset size increases.
@@
Our key contributions are as follows: (1) we introduce FEText, the first dataset of facial expression-related image-text pairs, providing a crucial resource for advancing FER tasks. (2) we propose SynFER, the first diffusion-based data synthesis pipeline for FER, integrating FAUs information and semantic guidance to achieve fine-grained control and faithful expression generation. (3) we develop FERAnno, a novel diffusion-based label calibrator, designed to automatically refine and enhance the annotations of synthesized facial expression images. (4) extensive experiments across various datasets and learning paradigms demonstrate the effectiveness of the proposed SynFER, validating the quality and scalability of its synthesized FER data.
@@ 2 Related Work
Facial Expression Recognition (FER): Recent success in deep learning (DL) has largely boosted the performance of the FER task, despite the substantial data requirements for training DL models. To address the limited training data in FER, previous methods mainly focus on developing different learning paradigms, including semi-supervised learning Li et al. (2022b); Yu et al. (2023); Cho et al. (2024), transfer learning Li et al. (2022c); Ruan et al. (2022) and multi-task learning Liu et al. (2023b); Li et al. (2023a). For example, Ada-CM Li et al. (2022b) learns a confidence margin to make full use of the unlabeled facial expression data in a semi-supervised manner. Despite achieving performance gains for FER, these methods remain constrained by limited data. Recently, researchers have explored an alternative data-driven perspective of introducing large-scale face datasets from other facial analysis tasks (e.g., face recognition Zeng et al. (2022)). Meta-Face2Exp Zeng et al. (2022) utilizes large-scale face recognition data to enhance FER by matching the feature distribution between face recognition and FER. However, face data drawn from these datasets lack diverse facial expressions, and thereby couldn’t fully unlock the potential of large-scale data in FER.
Synthetic Data: Recently, growing attention has been paid to the advanced generative models (e.g., Generative Adversarial Networks (GANs) Goodfellow et al. (2020) and Diffusion Models Rombach et al. (2022b)), which are typically flexible to synthesize training images for a wider range of downstream tasks, including classification Frid-Adar et al. (2018); Azizi et al. (2023), face recognition Kim et al. (2023); Boutros et al. (2023), semantic segmentation Nguyen et al. (2023); Wu et al. (2024; 2023a) and human pose estimation Feng et al. (2023); Zhou et al. (2023). In particular, some studies pioneer to investigate the capabilities of powerful pre-trained diffusion generative models on natural images Nguyen et al. (2023); Wu et al. (2024); Li et al. (2023b). For example, DatasetDM Wu et al. (2024) further introduces a generalized perception decoder to parse the rich latent space of the pre-trained diffusion model for various downstream tasks. However, there remains a significant gap in research on using diffusion models to generate facial expression data, including both images and corresponding labels. In this paper, we address this gap by exploring diffusion-based synthetic data for the first time in FER. Specifically, we propose to apply a fine-tuned diffusion model to facial expression synthesis and introduce the first diffusion-based pseudo-label generator for FER.
@@ 3 Preliminaries
Diffusion models include a forward process that adds Gaussian noise $\epsilon$ to convert a clean sample $x_{0}$ to noise sample $x_{T}$, and a backward process that iteratively performs denoising from $x_{T}$ to $x_{0}$, where $T$ represents the total number of timesteps. The forward process of injecting noise can be formulated as:
@@
$x_{t}$ is the noise feature at timestep $t$ and $\alpha_{t}$ is a predetermined hyperparameter for sampling $x_{t}$ with a given noise scheduler Song et al. (2020). In the backward process of denoising, given input noise $x_{t}$ sampled from a Gaussian distribution, a learnable network $\epsilon_{\theta}$ estimates the noise at each timestep $t$ with condition $c$. $x_{t-1}$, the feature at the previous timestep, is then derived as:
@@ 4 Methodology
We begin by introducing i) the overall synthetic pipeline for generating facial expression image-label pairs. Next, we detail ii) our approach for producing high-fidelity facial expression images, which are controlled through high-level text descriptions (Sec.4.2.1), fine-grained facial action units corresponding to localized facial muscles (Sec.4.2.3), and a semantic guidance technique (Sec.4.3). Finally, we introduce iii) the FER annotation crafter (FERAnno), a crucial component that thoroughly understands the synthetic facial expression data and automatically generates accurate annotations accordingly(Sec.4.4). This pipeline ensures both precision and reliability in facial expression generation and labeling.
@@ 4.1 Overall Pipeline for FER Data Synthesis
We introduce the overall pipeline for FER data synthesis (Fig. 2). The process starts with a coarse human portrait description assigned to a specific facial expression. ChatGPT enriches this description with details such as facial appearance, subtle facial muscle movements, and contextual cues. Simultaneously, facial action unit annotations are generated based on prior FAU-FE knowledge Ekman & Friesen (1978), aligning them with emotion categories to serve as explicit control signals for guiding the facial expression image synthesis. Once the facial expression label, facial action unit labels, and expanded textual prompt are prepared, these inputs condition our diffusion model to generate high-fidelity FER images, guided by semantic guidance to ensure accurate FER semantic. During the denoising process, FERAnno automatically produces pseudo labels for the generated images. To further improve labeling accuracy, we ensemble our FERAnno with existing FER models, which collaborate to vote on the accuracy of the predefined FER labels. In cases where discrepancies arise, the predefined label is refined by averaging the predictions from the ensemble experts. This mechanism effectively reduces the risk of inconsistent or uncertain annotations, ensuring that the final synthesis data is precise and dependable for downstream applications.
@@ 4.2.1 FEText Data Construction
To address the lack of facial expression image-text pairs for diffusion model training, we introduce FEText (Fig. 3), the first hybrid image-text dataset for FER. It combines face images from FFHQ Karras et al. (2019), CelebA-HQ Karras et al. (2017), AffectNet Mollahosseini et al. (2017) and SFEW Dhall et al. (2011), each paired with captions generated by a multi-modal large language model (MLLM). FEText includes 400,000 curated pairs tailored for facial expression tasks.
Resolution Alignment. Due to variations in image resolution across different datasets, we first utilize a super-resolution model Lin et al. (2023) to standardize the resolutions of images from AffectNet and SFEW. Specifically, we incorporate high-resolution images from FFHQ and CelebA-HQ datasets to preserve the model’s capacity for high-fidelity image generation. This dual approach allows the model to not only maintain the fidelity of the generated images but also to learn and incorporate the facial expression semantics from AffectNet and SFEW.
Textual Caption Annotation. To generate a textural caption for each face image, we employ the open-source multi-modal language model ShareGPT-4V Chen et al. (2023b), by guiding it with carefully crafted instructions. To ensure that the generated captions are both context-aware and expressive, we clearly define the model’s role and provide examples of detailed facial expression descriptions within the prompts. This approach enables the model to generate precise, emotion-reflective captions for the input images.
4.2.2 Diffusion Model Fine-tuning
To facilitate our diffusion model to generate high-fidelity facial expressions, a straightforward approach is to fine-tune the model directly on the proposed FEText using the diffusion loss in Eq.  3. However, since FEText contains images processed through a super-resolution model, this direct fine-tuning strategy may lead to over-smoothing in the generated images. To address this, we introduce a two-stage fine-tuning paradigm. In the first stage, the diffusion model is trained on the entire FEText dataset to capture facial expression-related semantics. Then, the second stage mitigates over-smoothing by specifically fine-tuning our diffusion model on the CelebA-HQ and FFHQ subsets of FEText, which consist of native high-resolution images. This two-step approach ensures that our model learns expressive facial details while preserving image sharpness. The fine-tuned model then serves as the foundation for controllable facial expression generation, incorporating facial action unit injection (Sec. 4.2.3) and semantic guidance (Sec. 4.3).
@@ 4.2.3 Explicit Control Signals via Facial Action Units
While fine-tuning the diffusion model using facial expression captions provides general language-based guidance for facial expression generation, it lacks the precision needed to capture fine-grained facial details, such as localized muscle movements. To overcome this limitation, we propose to incorporate more explicit control signals through Facial Action Units (FAUs), each of which represents a specific facial muscle movement. Inspired by IP-Adapter Ye et al. (2023), we apply a decoupled cross-attention module to integrate FAU embeddings with the diffusion model’s generation process. These embeddings are derived by mapping discrete FAU labels into high-dimensional representations using a Multi-Layer Perceptron, referred to as the AU adapter. FAU labels for each image in the FEText dataset are annotated using the widely adopted FAU detection model, OpenGraphAU Luo et al. (2022). With the diffusion model’s parameters frozen, we train the AU adapter to guide the model in recovering facial images based on the annotated FAU labels, using the objective in Eq. 3.

### v3

@@ Abstract
Facial expression datasets remain limited in scale due to the subjectivity of annotations and the labor-intensive nature of data collection. This limitation poses a significant challenge for developing modern deep learning-based facial expression analysis models, particularly foundation models, that rely on large-scale data for optimal performance. To tackle the overarching and complex challenge, instead of introducing a new large-scale dataset, we introduce SynFER (Synthesis of Facial Expressions with Refined Control), a novel synthetic framework for synthesizing facial expression image data based on high-level textual descriptions as well as more fine-grained and precise control through facial action units. To ensure the quality and reliability of the synthetic data, we propose a semantic guidance technique to steer the generation process and a pseudo-label generator to help rectify the facial expression labels for the synthetic images. To demonstrate the generation fidelity and the effectiveness of the synthetic data from SynFER, we conduct extensive experiments on representation learning using both synthetic data and real-world data. Results validate the efficacy of our approach and the synthetic data. Notably, our approach achieves a 67.23% classification accuracy on AffectNet when training solely with synthetic data equivalent to the AffectNet training set size, which increases to 69.84% when scaling up to five times the original size. Code is available here.
footnotetext: * These authors contributed equally.footnotetext: $\dagger$ indicates corresponding authors.
@@ 1 Introduction
Facial Expression Recognition (FER) is at the forefront of advancing AI’s ability to interpret human emotions, opening new frontiers for various human-centered applications. From automatic emotion detection to early interventions in mental health [63], accurate pain assessment [33], and enhancing human-computer interaction [1], the potential impact of FER systems is profound [57, 67, 91]. In recent years, learning-based FER models have gained significant traction due to their promising performances [42, 86, 24]. However, despite recent advancements in network architectures and learning methodologies, the progress of existing FER models has been hindered by the inadequate scale and quality of available training data, underscoring the need to expand datasets with high-quality data to push the boundaries of FER capabilities.
@@
Existing FER datasets, such as CK+ (953 sequences) [53], FER-2013 (30,000 $\times48$ images) [5], RAF-DB (29,672 images) [43], AFEW (113,355 images) [19], and SFEW (1,766 images) [18], are small compared to popular image datasets for general image processing (e.g., ImageNet [17] with 1.4 million images and Laion [68] with billion-level data). While AffectNet [58] compiles a large number of facial images from the web, it still suffers from vital drawbacks. A considerable portion of AffectNet’s images are low-quality, and its annotations often contain incorrect labels, which impairs the training process of FER models [38, 78]. Consequently, the absence of high-quality and large-scale FER datasets has delayed the development of FER foundation models. However, collecting a large-scale FER dataset with high-quality facial images and meticulous annotations is almost an unrealistic endeavor due to substantial financial and time costs, ethical concerns around facial data collection, and limited resources for large-scale acquisition. Additionally, the subjective interpretation of facial expressions results in inconsistent labeling by annotators, which exacerbates variability and hinders the creation of reliable datasets.
@@
To address the challenges in developing FER models, we turn to explore the generation paradigm to synthesize high-quality facial expression images paired with reliable labels. This approach draws inspiration from successful strategies employed to expand annotated datasets for other computer vision tasks, such as semantic segmentation [4, 14, 39] and depth estimation [2, 15, 30]. These advances leverage powerful generative models such as Stable Diffusion [64] and DALL-E [6], which capture intricate natural image patterns. By tapping into these models, researchers have generated realistic images with their corresponding annotations, thereby boosting model performance. However, applying diffusion models to synthesize facial expression images with reliable FER labels presents two major challenges. (1) the training sets used by these generative models often lack diverse facial expression data, limiting their ability to produce images that capture subtle and nuanced emotional semantics; and (2) prior approaches to generate annotations for synthetic images focused on tangible attributes such as pixel-wise layouts, or depth maps. In contrast, facial expressions convey abstract and subjective emotions, making the generation of precise and reliable expression labels much more complex. To the best of our knowledge, none of the existing methods can simultaneously conduct fine-grained control for facial expression generation and generate robust categorical facial expression labels leveraging diffusion-based features.
@@
In this paper, instead of a dataset, we present SynFER, a leading data synthesis pipeline capable of synthesizing unlimited and realistic facial expression images paired with reliable expression labels, to drive advancements in FER models. To address the shortcomings of existing FER datasets, which often lack expression-related text paired with facial images, we introduce FEText, a unique hybrid dataset created by curating and filtering data from existing FER and high-quality face datasets. This vision-language dataset serves as the foundation for training our generative model to synthesize facial expression data. To ensure fine-grained control and faithful generation of facial expression images, we inject facial action unit (FAU) information and semantic guidance from external pre-trained FER models. Building upon this, we propose FERAnno, the first diffusion-based label calibrator for FER, which automatically generates reliable annotations for the synthesized images. Together, these innovations position SynFER as a powerful tool for producing large-scale, high-quality facial expression data, offering a significant resource for the development of FER models.
@@
We investigate the effectiveness of the synthetic data across various learning paradigms, demonstrating consistent and modest improvement in model performance. As shown in Fig. 1(c), training with the synthetic data yields significant performance boosts across various learning paradigms. Notably, pre-training on the synthetic data (Fig. 1(c)) improves self-supervised learning model MoCov3 [13] on AffectNet, surpassing real-world data pre-training. In supervised learning, SynFER improves accuracy by +1.55% for the state-of-the-art FER model, POSTER++ [56], on AffectNet. We further explore the performance scaling of the synthetic data, revealing further gains as dataset size increases. Our key contributions are:
  • We introduce FEText, the first dataset of facial expression-related image-text pairs, providing a crucial resource for advancing FER tasks.
@@
  • We propose SynFER, the first diffusion-based data synthesis pipeline for FER, integrating FAUs information and semantic guidance to achieve fine-grained control and faithful expression generation. Additionally, FERAnno, a novel diffusion-based label calibrator, is designed to automatically refine and enhance the annotations of synthesized facial expression images.
  • Extensive experiments across six datasets and four learning paradigms demonstrate the effectiveness of the proposed SynFER, validating the quality and scalability of its synthesized output.
@@ 2 Related Work
Facial Expression Recognition (FER): Recent success in deep learning (DL) has largely boosted the performance of the FER task, despite the substantial data requirements for training DL models. To address the limited training data in FER, previous methods mainly focus on developing different learning paradigms, including semi-supervised learning [41, 81, 16], transfer learning [45, 66] and multi-task learning [52, 44]. For example, Ada-CM [41] learns a confidence margin to make full use of the unlabeled facial expression data in a semi-supervised manner. Despite achieving performance gains for FER, these methods remain constrained by limited data. Recently, researchers have explored an alternative data-driven perspective of introducing large-scale face datasets from other facial analysis tasks (e.g., face recognition [84]). Meta-Face2Exp [84] utilizes large-scale face recognition data to enhance FER by matching the feature distribution between face recognition and FER. However, face data drawn from these datasets lack diverse facial expressions, and thereby couldn’t fully unlock the potential of large-scale data in FER.
Synthetic Data: Recently, growing attention has been paid to the advanced generative models (e.g., Generative Adversarial Networks (GANs) [28] and Diffusion Models [65]), which are typically flexible to synthesize training images for a wider range of downstream tasks, including classification [26, 3], face recognition [37, 7], semantic segmentation [60, 75, 74] and human pose estimation [25, 90]. In particular, some studies pioneer the capabilities of powerful pre-trained diffusion generative models on natural images [60, 75, 46]. For example, DatasetDM [75] further introduces a generalized perception decoder to parse the rich latent space of the pre-trained diffusion model for various downstream tasks. Despite the growing adoption of diffusion models in synthetic data generation, their application to anatomically grounded facial expression synthesis remains critically underexplored. While recent diffusion-based methods like FineFace [72] and EmojiDiff [34] achieve visually compelling facial expression synthesis through AU-guided editing or reference-driven generation, their utility as synthetic training data for FER remains unverified. However, these methods prioritize perceptual quality over functional utility, lacking explicit mechanisms to align generated expressions with FER label semantics or to validate downstream task performance. In this paper, we attempt to investigate the potential and feasibility of synthetic images for FER tasks.
@@ 3 Preliminaries
Diffusion models include a forward process that adds Gaussian noise $\epsilon$ to convert a clean sample $x_{0}$ to noise sample $x_{T}$, and a backward process that iteratively performs denoising from $x_{T}$ to $x_{0}$, where $T$ represents the total number of timesteps. The forward process of injecting noise is:
@@
$x_{t}$ is the noise feature at timestep $t$ and $\alpha_{t}$ is a predetermined hyperparam. for sampling $x_{t}$ with a given noise scheduler [69]. In backward process of denoising, given input noise $x_{t}$ sampled from a Gaussian distribution, a learnable network $\epsilon_{\theta}$ estimates the noise at each timestep $t$ with condition $c$. $x_{t-1}$, the feature at the previous timestep is:
@@ 4 Methodology
We first introduce i) the overall synthetic pipeline for generating facial expression image-label pairs. Next, we detail ii) our approach for producing high-fidelity facial expression images, which are controlled through high-level text descriptions (Sec.4.2.1), fine-grained facial action units corresponding to localized facial muscles (Sec.4.2.2), and a semantic guidance technique (Sec.4.3). Finally, we introduce iii) the FER annotation crafter (FERAnno), a crucial component that thoroughly understands the synthetic facial expression data and automatically generates accurate annotations accordingly (Sec.4.4). This pipeline ensures both precision and reliability in facial expression generation and labeling.
@@ 4.1 Overall Pipeline for FER Data Synthesis
We introduce the overall pipeline for FER data synthesis (Fig. 2). The process starts with a coarse human portrait description assigned to a specific facial expression. ChatGPT enriches this description with details such as facial appearance, subtle facial muscle movements, and contextual cues. Simultaneously, facial action unit annotations are generated based on prior FAU-FE knowledge [21], aligning them with emotion categories to serve as explicit control signals for guiding the facial expression image synthesis. Once the facial expression label, facial action unit labels, and expanded textual prompt are prepared, these inputs condition our diffusion model to generate high-fidelity FER images, guided by semantic guidance to ensure accurate FER semantics. During the denoising process, FERAnno automatically produces pseudo labels for the generated images. To further improve labeling accuracy, we ensemble our FERAnno with existing FER models, which collaborate to vote on the accuracy of the predefined FER labels. In cases where discrepancies arise, the predefined label is replaced by averaging the predictions from the ensemble experts. This mechanism effectively reduces the risk of inconsistent or uncertain annotations, ensuring that the final synthesis data is precise and dependable for downstream applications.
@@ 4.2.1 FEText Data Construction
To address the lack of facial expression image-text pairs for diffusion model training, we introduce FEText (Fig. 3), the first hybrid image-text dataset for FER. It combines face images from FFHQ [36], CelebA-HQ [35], AffectNet [58], and SFEW [18], each paired with captions generated by a multi-modal large language model (MLLM). FEText includes 400K curated pairs tailored for facial expression tasks.
Resolution Alignment: Due to variations in image resolution across different datasets, we first utilize a super-resolution model [48] to standardize the resolutions of images from AffectNet and SFEW. Specifically, we incorporate high-resolution images from FFHQ and CelebA-HQ datasets to preserve the model’s capacity for high-fidelity image generation. This dual approach allows the model to not only maintain the fidelity of the generated images but also to learn and incorporate the facial expression semantics from AffectNet and SFEW.
Textual Caption Annotation: To generate a textural caption for each face image, we employ the open-source multi-modal language model ShareGPT-4V [11], by guiding it with carefully crafted instructions. To ensure that the generated captions are both context-aware and expressive, we clearly define the model’s role and provide examples of detailed facial expression descriptions within the prompts. This approach enables the model to generate precise, emotion-reflective captions for the input images.
With the FEText obtained, we then conduct the fine-tuning of the diffusion model on it using the diffusion loss in Eq. 3. A detailed fine-tuning strategy of the model can be found in the supplementary material.
@@ 4.2.2 Explicit Control Signals via Facial Action Units
While fine-tuning the diffusion model using facial expression captions provides general language-based guidance for facial expression generation, it lacks the precision needed to capture fine-grained facial details, such as localized muscle movements. To this end, we propose to incorporate more explicit control signals through Facial Action Units (FAUs), each of which represents a specific facial muscle movement. Inspired by IP-Adapter [80], we apply a decoupled cross-attention module to integrate FAU embeddings with the diffusion model’s generation process. These embeddings are derived by mapping discrete FAU labels into high-dimensional representations using a Multi-Layer Perceptron, referred to as the AU adapter. FAU labels for each image in the FEText dataset are annotated using the widely adopted FAU detection model, OpenGraphAU [54]. With the diffusion model’s parameters frozen, we train the AU adapter to guide the model in recovering facial images based on the annotated FAU labels, using the objective in Eq. 3.